OpenStudy (anonymous):

1/((sqrtx)(sqrtx+1)) dx integral

OpenStudy (anonymous):

is it (sqrtx +1) or sqrt(x+1)?

OpenStudy (anonymous):

\[1/((\sqrt{x})(\sqrt{x}+1))\]

OpenStudy (anonymous):

the answer from the book is\[2\ln \left| \sqrt{x}+1 \right|+c\]

OpenStudy (anonymous):

this is equal to dx/(x + sqrtx) . let u = sqrtx so du = 1/2 (x^(-1/2)) * dx this is equal to [1/(u^2 + u)] * 2u du = 2u/(u^2 + u) lmk if you would like the next step

OpenStudy (anonymous):

ok i substituted sqrtx back in but its not the book answer

OpenStudy (anonymous):

ooh the book say use (sqrtx+1)=u

OpenStudy (anonymous):

2u/(u^2 + u) = 2/(u +1) --> simpler to integrate

OpenStudy (anonymous):

both substitutions work. in fact even a whacky substitution like u = sqrtx + pi could work if done right. integral is 2 ln |u + 1| + C . u = sqrtx. ^_^

OpenStudy (anonymous):

thanks you so much for helping me